# BanTH: A Multi-label Hate Speech Detection Dataset for Transliterated Bangla

### v3

@@ Abstract
The proliferation of transliterated texts in digital spaces has emphasized the need for detecting and classifying hate speech in languages beyond English, particularly in low-resource languages. As online discourse can perpetuate discrimination based on target groups, e.g. gender, religion, and origin, multi-label classification of hateful content can help in comprehending hate motivation and enhance content moderation. While previous efforts have focused on monolingual or binary hate classification tasks, no work has yet addressed the challenge of multi-label hate speech classification in transliterated Bangla. We introduce BanTH, the first multi-label transliterated Bangla hate speech dataset comprising 37.3k samples. The samples are sourced from YouTube comments, where each instance is labeled with one or more target groups, reflecting the regional demographic. We establish novel transformer encoder-based baselines by further pre-training on transliterated Bangla corpus. We also propose a novel translation-based LLM prompting strategy for transliterated text. Experiments reveal our further pre-trained encoders achieving state-of-the-art performance on the BanTH dataset while our translation-based prompting outperforms other strategies in the zero-shot setting. The introduction of BanTH not only fills a critical gap in hate speech research for Bangla but also sets the stage for future exploration into code-mixed and multi-label classification challenges in underrepresented languages.
@@
Note: Throughout the work, we use the term Bangla to refer to both Bangla and the endonym Bengali, both denoting the same language primarily spoken by people of the West Bengal region of India and the vast majority of Bangladesh.
@@ 1 Introduction
The ever-expanding digital landscape that promises to improve social cohesion has been a breeding ground for hate speech. Hate speech is defined as any form of language that targets, attacks, or incites implicit or explicit forms of hatred or violence against groups, based on specific characteristics, e.g. physical appearance, religion, national or ethnic origin, and gender identity. Hate speech can potentially inflict societal harm by promoting division among communities, exacerbating mental health problems, and inciting violence. The lack of hate speech moderation cultivates an environment of intolerance and magnifies the negative impact on the target communities Hangartner et al. (2021). Furthermore, categorizing hate speech requires a multi-faceted approach to capture overlapping hate categories and gain a granular understanding of the underlying motives behind the hateful discourse.
A prevalent informal form of online communication for non-English languages uses transliterated texts, i.e. writing one language in the script of another. Transliteration is predominant in Bangla, where Latin characters are used to produce a colloquial form of Bangla texts that does not strictly adhere to the original linguistic rules. Despite boasting over a quarter of a billion speakers worldwide, Bangla remains a low-resource language in terms of Natural Language Processing (NLP) tasks. The challenges are compounded when working with the even more under-resourced transliterated Bangla, which dominates online spaces due to the widespread familiarity of users with English keyboard layouts.
Although transliteration facilitates easier typing and cross-linguistic communication, it also complicates the task of automated hate speech detection due to inconsistent spelling and structure, absence of grammar, mixing with English, and the loss of script-specific features Jahan et al. (2019a). While standard practices have employed transformer-based encoders Devlin et al. (2019) for automated hate speech detection, the recent surge in popularity of Large Language Models (LLMs) has positioned them as a viable option in hate speech NLP, particularly in the zero-shot setting. While most of the advancements have been centered on English and other high-resource languages, transliterated Bangla has very limited research on hate or hate-like speech detection Jahan et al. (2019b); Raihan et al. (2023) and LLM-based methods Shibli et al. (2023), with no work addressing the intersection of these areas.
@@
We establish several encoder-based baselines along with our novel transliterated Bangla further-pretrained encoders, achieving state-of-the-art accuracy on our dataset.
@@
2 BanTH Dataset
The BanTH dataset comprises 37350 samples, each initially classified into binary labels of Hate or Non-Hate. To effectively capture the target group, the hate samples are further multi-labeled as Political, Religious, Gender, Personal Offense, Abusive/Violence, Origin, and Body Shaming. A detailed description of the target group is provided in Appendix C.1. The flowchart of the dataset creation is given in Fig. 2.
2.1 Data Sourcing
We construct the dataset by scraping user comments from public YouTube videos using the YouTube API111https://developers.google.com/youtube/v3/getting-started. The scraped comments include three categories of videos: “News & Politics”, “People & Blogs”, and “Entertainment”, totaling 26 different YouTube channels, outlined in Table 1. The contents often cover a wide range of topics, including political analysis, interviews, talk shows, product reviews, and travel videos, regionally relevant to West Bengal of India and Bangladesh.
The scraped comments were collected from January 2020 to July 2024, a time frame that captures significant recent events, e.g. the COVID-19 pandemic and the Bangladesh quota reform movement222https://en.wikipedia.org/wiki/2024_Bangladesh_quota_reform_movement. Such events led to increased social media activity. The variety of content appeals to a diverse audience with a wide range of interests and perspectives. The label-wise distribution of video categories is elaborated in Appendix C.2.
2.2 Data Filtering
We primarily filtered comments containing English characters. During annotation, non-Bangla and non-English transliterated comments, such as comments in Hindi, were discarded. The scraped comments were filtered through English character filtering and were further filtered using dictionary-based language identification. Original English and Bangla and code-mixed texts were filtered out using the dictionary filtering approach. Transliterated texts in any other language such as Hindi, were dropped during annotation. Only the transliterated Bangla texts with lengths greater than or equal to three were kept in the dataset.
2.3 Data Cleaning
During data cleaning, we removed URLs using regular expressions as they were not relevant to our datasets. Comments that were verbatim duplicates of others were removed. Personal information like addresses and contact numbers were also removed during annotation.
2.4 Data Annotation
We hired four annotators and two domain experts for the data annotation process. All the annotators were high school graduates and native Bangla speakers. They had extensive exposure to social media content and actively used transliterated Bangla text. The annotators were compensated fairly for their work, with a rate of BDT 1 per sample. The domain experts were remunerated hourly. The annotation process was supervised by domain experts with experience in working with transliterated Bangla text, who provided guidance and ensured clear and proper annotation guidelines, outlined in Appendix B.
Each data sample was annotated by three annotators to ensure consistency and capture a shared perspective. Each data sample was first instructed to be classified as hate or non-hate and each resulting hate sample was then instructed to be multi-labeled according to the target group descriptions given above. In cases where annotators encountered difficulties in reaching a consensus or understanding the text, these issues were addressed and instantly resolved by the domain experts to avoid further confusion.
| Fleiss’ Kappa | Expert- Annotator | Inter- Annotator |
| Binary Labeling | 0.75 | 0.71 |
| Political | 0.62 | 0.72 |
| Religious | 0.68 | 0.66 |
| Gender | 0.64 | 0.68 |
| Personal Offense | 0.70 | 0.75 |
| Abusive/Violence | 0.72 | 0.71 |
| Origin | 0.68 | 0.72 |
| Body Shaming | 0.64 | 0.66 |
2.5 Data Validation
Following the annotation process, the dataset underwent a thorough review by the domain experts to ensure labeling accuracy. The experts verified whether any samples remained unlabeled, whether any samples were mislabeled, or if there were any instances where a sample was labeled as hate without the corresponding multi-label, and vice versa. Additionally, the dataset was examined by the domain experts to ensure that all samples were in transliterated Bangla, as samples in other languages were considered beyond the scope of the study and were subsequently removed. The agreement between the annotators the domain experts, and the inter-annotator was measured using the Fleiss’ Kappa metric and the result is shown the Table 2. An inter-annotator agreement score of 0.53 indicates a moderate agreement across the dataset Islam et al. (2021).
@@ 2.6 Dataset Statistics
Following data validation, the dataset is split into train, test, and val datasets maintaining the standard 80:10:10 ratio. During the split, a stratification approach was followed to maintain a uniform distribution of hate and non-hate labels and further multi-label categories across all the splits. Further data statistics are given in Table 1.
The heatmap in Figure 3 provides a detailed visualization of multi-label co-occurrence within the BanTH dataset, illustrating the frequency with which pairs of labels appear together in the text data. Each cell in the heatmap represents the count of co-occurrence between two specific labels, with darker shades of purple indicating higher levels of co-occurrence. The diagonal cells depict the total occurrences of each label independently, with Personal Offense emerging as the most frequent category, appearing in 6966 samples. Conversely, Body Shaming is the least frequent label in the dataset. Significant co-occurrence patterns are observed between Personal Offense and Abusive/Violence with 618 samples, as well as between Personal Offense and Political with 472 samples. In contrast, labels such as Religious and Body Shaming demonstrate minimal co-occurrence with other categories, reflecting their relative isolation in the dataset. The heatmap underscores the prominence of Personal Offense in the BanTH dataset, indicating that a substantial portion of the text data is classified under this label.
@@ 3 Methodology
We use various baseline approaches to assess the quality of BanTH, establishing the benchmark performance of it. The experiments are carried out in two stages: Binary Classification of the dataset, followed by Multi-label classification of the hate speech predicted in the first stage. For evaluation, we follow two approaches: one involves finetuning the pre-trained LMs on our corpus, BanTH, and the other involves prompting state-of-the-art large GPT-based LLMs. The following sections discuss the experiments in detail.
3.1 Further Pretraining
Pretrained LMs such as BERT Devlin et al. (2019) and BanglaBERT Bhattacharjee et al. (2022) are typically trained on general and formal texts. However, these models often lack transliterated samples in their pretraining datasets. To enhance the performance of these models, we employ the technique of Further Pretraining (FPT) Qiu et al. (2021); Gururangan et al. (2020), utilizing the Masked Language Modeling (MLM) loss Devlin et al. (2019); Zhuang et al. (2021) as our pretraining objective. This approach aims to improve the models’ effectiveness on our BanTH dataset.
For our FPT process, we utilize the BanglaTLit-PT corpus333https://www.kaggle.com/datasets/farihatanjimshifat1/bangla-transliteration-further-pretraining-dataset, which contains 243k unlabeled Bangla transliterated texts. After further pretraining the models on BanglaTLit-PT, we build Bangla transliteration-enhanced encoder models namely TB-Encoders (Transliterated Bangla Encoders).
3.2 LM Finetuning
In our fine-tuning experiments, we utilize a pre-trained language model that has been specifically trained on our dataset. We input a sentence into to obtain layer-wise contextual representations , where and = no. of layers in . We use the representation from the last layer’s [CLS] token, , as the sentence representation, which is then passed through a Multi-Layer Perceptron (MLP) for classification. The transformation is defined as follows:
The resulting representation is then employed for calculating the loss. In the case of TB-Encoders, we also similarly fine-tune those models after doing FPT on the BanglaTLit-PT dataset.
@@ 3.3 Prompting Strategy
We adopt one of the following strategies: Why Positive/Why Negative Wang et al. (2023), HARE Yang et al. (2023), non-explanatory, explanation-based, CoT based or translation based. An overview of the prompts has been reported in Tab. 3. We consider GPT 4o and GPT-3.5 in the prompting experiments and have designed separate base prompts for binary and multi-label classification. Each of these base prompts is then extended to one of the prompting strategies. We consider both zero-shot and few-shot prompting techniques, reported in App. E. Overall, all the variations of the prompts can be categorized as follows:
Non-Explanatory. The non-explanatory prompting strategy forms the base prompts for other prompting strategies, where we only ask the LLM to do binary or multi-label classification. The prompt also includes the hate speech definition, the target groups, and their definition in multi-label classification, key indicators of hate, and labeling instructions.
Explanation Based. In the explanation-based prompting approach, we include the question Explain why with the base prompt and ask the LLM to return its reasoning.
HARE. In HARE-based prompting strategy, we insert the line Let’s explain step by step, as adapted in Yang et al. (2023), along with the base-prompt and asked the LLM to return its reason for the corresponding binary or multi-label.
Why Positive Why Negative. As explained in Wang et al. (2023), we adopt the approach by adding the question why the given text is positive or negative appending to the base prompt in this variation.
CoT-Based. CoT-based prompting strategy is implemented by adding the question Let’s think step by step along with the base prompt.
Translation Based. LLMs are trained primarily on extensive formal text, which may limit their performance and explanatory capabilities when handling transliterated text. To address this issue, we propose a simple yet effective prompting strategy called the Translation-Based prompting strategy. This approach requires LLMs to first convert Bengali transliterated text into standard Bangla or English before proceeding with the usual classification. We test this proposed prompting method using Non-Explanatory, Explanatory, and CoT-based prompting techniques.
All the prompts are provided in Appendix E.
@@
| BanglaBERT | 76.50 | 81.04 | 20.82 | 54.08 | 7.26 |
| BanglishBERT | 75.07 | 80.62 | 20.61 | 52.74 | 7.42 |
| BanglaHateBERT | 70.92 | 77.54 | 11.34 | 49.83 | 7.97 |
| VAC-BERT | 74.19 | 79.76 | 18.45 | 52.56 | 7.50 |
@@
| MuRIL | 75.29 | 80.83 | 14.58 | 53.98 | 7.55 |
| IndicBERT | 74.51 | 80.48 | 13.39 | 51.56 | 7.88 |
@@
| CharBERT | 76.61 | 80.91 | 19.66 | 53.21 | 7.44 |
| Transliterated Bangla (TB) LM (Ours) | |||||
@@
| Non-Explanatory | |||||
| GPT 3.5 | 61.44 | 64.02 | 24.12 | 18.27 | 19.77 |
| GPT 4o | 70.05 | 74.30 | 36.07 | 22.60 | 16.91 |

### v4

@@ Abstract
The proliferation of transliterated texts in digital spaces has emphasized the need for detecting and classifying hate speech in languages beyond English, particularly in low-resource languages. As online discourse can perpetuate discrimination based on target groups, e.g. gender, religion, and origin, multi-label classification of hateful content can help in understanding hate motivation and enhance content moderation. While previous efforts have focused on monolingual or binary hate classification tasks, no work has yet addressed the challenge of multi-label hate speech classification in transliterated Bangla. We introduce BanTH, the first multi-label transliterated Bangla hate speech dataset. The samples are sourced from YouTube comments, where each instance is labeled with one or more target groups, reflecting the regional demographic. We propose a novel translation-based LLM prompting strategy that translates or transliterates under-resourced text to higher-resourced text before classifying the hate group(s). Experiments reveal further pre-trained encoders achieving state-of-the-art performance on the BanTH dataset while translation-based prompting outperforms other strategies in the zero-shot setting. We address a critical gap in Bangla hate speech and set the stage for further exploration into code-mixed and multi-label classification in underrepresented languages.
@@
Note: Throughout the work, we use the term Bangla referring to both Bangla and the endonym Bengali. The terms denote the same language, primarily spoken by people of the West Bengal region of India and the vast majority of Bangladesh.
@@ 1 Introduction
The ever-expanding digital landscape that promises to improve social cohesion has been a breeding ground for hate speech Castaño-Pulgarín et al. (2021). Hate speech is defined as any form of language that targets, attacks, or incites implicit or explicit forms of hatred or violence against groups, based on specific characteristics, e.g. physical appearance, religion, ethnic origin, and gender identity Papcunová et al. (2023). Hate speech can potentially inflict societal harm by promoting division among communities, exacerbating mental health problems, and inciting violence Sahoo et al. (2024). The lack of hate speech moderation cultivates an environment of intolerance and magnifies the negative impact on the target communities Hangartner et al. (2021). Categorizing hate speech requires a multi-faceted approach to capture overlapping hate categories and gain a granular understanding of the underlying motives behind the hateful discourse.
A prevalent informal form of online communication for non-English languages uses transliterated texts, i.e. writing a language in foreign script. Transliteration is predominant in Bangla, where Latin characters are used to produce a colloquial form of Bangla texts that does not strictly adhere to the original linguistic rules. Despite boasting over a quarter of a billion speakers worldwide, Bangla remains a low-resource language in terms of Natural Language Processing (NLP) tasks Mahfuz et al. (2025). The challenges are compounded when working with the even more under-resourced transliterated Bangla, which dominates online spaces due to the widespread familiarity of users with English keyboard layouts Fahim et al. (2024).
Although transliteration facilitates easier typing and cross-linguistic communication, it also complicates the task of automated hate speech detection due to inconsistent spelling and structure, absence of grammar, mixing with English, and the loss of script-specific features Jahan et al. (2019a). While standard practices have employed transformer-based encoders Devlin et al. (2019) for automated hate speech detection, the recent surge in popularity of Large Language Models (LLMs) has positioned them as a viable option in hate speech NLP, particularly in the zero-shot setting. While most of the advancements have been centered on English and other high-resource languages, transliterated Bangla has limited research on hate or hate-like speech detection Jahan et al. (2019b); Raihan et al. (2023) and LLM-based methods Shibli et al. (2023), with no work addressing the intersection of these areas.
@@
We establish several encoder-based baselines, with encoders further pre-trained on transliterated Bangla achieving state-of-the-art accuracy on our dataset.
@@
| Hate Category | #News & Politics | #People & Blogs | #Enter- tainment | #Train | #Val | #Test | #Total |
| Political | 1005 | 174 | 21 | 958 | 120 | 122 | 1200 |
| Religious | 109 | 30 | 0 | 111 | 14 | 14 | 139 |
| Gender | 151 | 38 | 4 | 154 | 19 | 20 | 193 |
| Personal Offense | 5883 | 784 | 297 | 5572 | 696 | 696 | 6964 |
| Abusive/Violence | 2491 | 276 | 121 | 2312 | 287 | 289 | 2888 |
| Origin | 201 | 11 | 2 | 171 | 21 | 22 | 214 |
| Body Shaming | 56 | 29 | 1 | 68 | 8 | 10 | 86 |
| Time Period | Event | #Samples | Major Hate Categories |
| Dec’19 - Dec’22 | COVID-19 | 878 | Abusive/Violence, Personal |
| Dec’22 | Woman dragged by car to death | 1245 | Personal, Abusive/Violence |
| Feb’24 | BCL leader rapes woman in JU | 1542 | Personal, Abusive/Violence, Political |
| July’24 - Aug’24 | Quota reform movement | 4522 | Political, Abusive/Violence, Personal |
| Aug’24 | Anti-Hindu Violence | 1081 | Religious, Abusive/Violence |
2 BanTH Dataset
The BanTH dataset comprises 37350 samples, each initially classified into binary labels of Hate or Non-Hate. To effectively capture the target group, the hate samples are further multi-labeled as Political, Religious, Gender, Personal Offense, Abusive/Violence, Origin, and Body Shaming. The description of target groups is provided in §C.1 and the dataset creation flowchart is given in Fig. 2.
2.1 Data Sourcing
We construct the dataset by scraping user comments from public YouTube videos using the YouTube API111https://developers.google.com/youtube/v3/getting-started. The scraped comments include three categories of videos: “News & Politics”, “People & Blogs”, and “Entertainment”, totaling 26 different YouTube channels, outlined in Tab. 4. The contents often cover a wider range of topics, including political analysis, interviews, talk shows, product reviews, and travel videos, regionally relevant to West Bengal of India and Bangladesh.
The scraped comments were collected from January 2020 to July 2024, a time frame that captures significant recent events, e.g. the COVID-19 pandemic and the Bangladesh quota reform movement222https://en.wikipedia.org/wiki/2024_Bangladesh_quota_reform_movement. Such events led to increased social media activity. The variety of content appeals to a diverse audience with a wide range of interests and perspectives. The label-wise distribution of video categories is elaborated in Tab. 1 and event-wise major hate categories are described in Tab. 2.
2.2 Data Filtering
We primarily filtered comments written in Roman scripts and discarded comments in other transliterated languages, e.g. Hindi, during annotation. The scraped comments were further filtered using dictionary-based language identification to ensure only text from the desired language remains. Finally, texts with a length greater than or equal to three were kept in the dataset.
2.3 Data Cleaning
During data cleaning, we removed URLs using regular expressions as they were irrelevant to our task. We also removed comments that were verbatim duplicates of others. Personal information like addresses and contact numbers was also removed during annotation.
2.4 Data Annotation
We hired four annotators and two domain experts for the data annotation process. All the annotators were high school graduates and native Bangla speakers. They had extensive exposure to social media content and actively used transliterated Bangla text. The annotation process was supervised by domain experts with experience in working with transliterated Bangla text, who provided guidance and ensured clear and proper annotation guidelines, outlined in Appendix B. Both the annotators and domain experts received monetary compensation.
Each data sample was annotated by three annotators to ensure consistency and capture a shared perspective. The annotators were first instructed to classify a sample as hate or non-hate. The resulting hate sample was then multi-labeled according to the target group descriptions given in §C.1. In cases where annotators encountered difficulties in reaching a consensus or understanding the text, the domain experts resolved the issues.
| Fleiss’ Kappa | Expert- Annotator | Inter- Annotator |
| Binary Labeling | 0.75 | 0.71 |
| Political | 0.62 | 0.72 |
| Religious | 0.68 | 0.66 |
| Gender | 0.64 | 0.68 |
| Personal Offense | 0.70 | 0.75 |
| Abusive/Violence | 0.72 | 0.71 |
| Origin | 0.68 | 0.72 |
| Body Shaming | 0.64 | 0.66 |
2.5 Data Validation
Following the initial phase of data annotation, the annotators verified whether any samples remained unlabeled, samples were mislabeled, or if there were instances where a sample was labeled as hate without the corresponding multi-label, and vice versa. Additionally, they ensured that all samples were in transliterated Bangla only, as samples in other languages are beyond the scope of our work. The agreement scores, measured using the Fleiss’ Kappa, are shown the Table 3. We found an inter-annotator agreement score of 0.53, indicating a moderate agreement across the hate categories Islam et al. (2021).
@@ 2.6 Dataset Statistics
The train, test, and val dataset splits maintain the standard 80:10:10 ratio. While splitting, stratification was followed to maintain a uniform distribution of hate and non-hate labels in multi-label categories. Tab. 1 shows the distribution across video categories and dataset splits while general statistics is given Tab. 4. Fig. 3 visualizes the multi-label co-occurrence within the BanTH dataset, illustrating the frequency of label pairs appearing together in the text. Hate categories like Personal and Political frequently co-occur with other categories while the opposite is true for categories like Religious and Body Shaming.
@@ 3 Methodology
The experiments are carried out in two stages: binary classification of hate speech and multi-label (target group) classification of the predicted hate speech. Our baselines can be divided into two approaches: fine-tuning LMs on our dataset and prompting GPT-based LLMs using different prompting and shot settings.
@@
3.1 Further Pretraining
Pretrained LMs such as BERT Devlin et al. (2019) and BanglaBERT Bhattacharjee et al. (2022) are typically trained on general and formal texts. However, these models often lack transliterated samples in their pretraining datasets, essentially becoming a long-tail distribution problem. To enhance the performance, we employ Further Pretraining (FPT) Qiu et al. (2021); Gururangan et al. (2020), utilizing the Masked Language Modeling (MLM) loss Devlin et al. (2019); Zhuang et al. (2021) as our pretraining objective.
We use the encoders further pre-trained on BanglaTLit-PT corpusFahim et al. (2024), which contains 243k unlabeled Bangla transliterated texts. The encoders are referred to as TB-Encoders (Transliterated Bangla Encoders).
3.2 LM Finetuning
We input a sentence into a pre-trained language model to obtain layer-wise contextual representations , where and = no. of layers in . We use the representation from the last layer’s [CLS] token, , as the sentence representation, which is then passed through a Multi-Layer Perceptron (MLP) for classification. The transformation is defined as follows:
The resulting representation is then employed for calculating the loss. The TB-Encoders are similarly fine-tuned.
@@ 3.3 Prompting Strategy
We adopt one of the following strategies: Why Positive/Why Negative Wang et al. (2023), HARE Yang et al. (2023), non-explanatory, explanation-based, CoT Wei et al. (2022) based or translation based. An overview of the prompts has been reported in Tab. 5. We consider GPT 4o Achiam et al. (2023) and GPT-3.5 Brown et al. (2020) in the prompting experiments and have designed separate base prompts for binary and multi-label classification. Each of these base prompts is then extended to one of the prompting strategies. We consider both zero-shot Radford et al. (2019) and few-shot Brown et al. (2020) prompting techniques, reported in App. F. The prompt variations are detailed below:
Vanilla/Base. The base prompting strategy where we only ask the LLM to do binary or multi-label classification. The prompt also includes the hate speech definition, the target groups, and their definition in multi-label classification, key indicators of hate, and labeling instructions.
Explanation Based. We include the phrase Explain why with the base prompt and ask the LLM to show its reasoning.
HARE. We add Let’s explain step by step, as adapted in Yang et al. (2023), along with the base prompt and ask the LLM to return its reasoning.
Why Positive Why Negative. Following Wang et al. (2023), we adopt the approach by adding the question why the given text is positive or negative appending to the base prompt in this variation.
CoT-Based. Following Wei et al. (2022), we add Let’s think step by step along with the base prompt.
@@
| BanglaBERT Bhattacharjee et al. (2022) | 76.50 | 81.04 | 20.82 | 54.08 | 7.26 |
| BanglishBERT Bhattacharjee et al. (2022) | 75.07 | 80.62 | 20.61 | 52.74 | 7.42 |
| BanglaHateBERT Jahan et al. (2022a) | 70.92 | 77.54 | 11.34 | 49.83 | 7.97 |
| VAC-BERT Bhattacharyya et al. (2023) | 74.19 | 79.76 | 18.45 | 52.56 | 7.50 |
@@
| MuRIL Khanuja et al. (2021) | 75.29 | 80.83 | 14.58 | 53.98 | 7.55 |
| IndicBERT Kakwani et al. (2020) | 74.51 | 80.48 | 13.39 | 51.56 | 7.88 |
@@
| CharBERT Ma et al. (2020) | 76.61 | 80.91 | 19.66 | 53.21 | 7.44 |
| Transliterated Bangla (TB) LM Fahim et al. (2024) | |||||
@@
| Vanilla/Base Prompting | |||||
| GPT 3.5 Brown et al. (2020) | 61.44 | 64.02 | 24.12 | 18.27 | 19.77 |
| GPT 4o Achiam et al. (2023) | 70.05 | 74.30 | 36.07 | 22.60 | 16.91 |
@@
Translation Based. Our proposed prompting strategy requires LLMs to first convert Bengali transliterated text into standard Bangla or English before proceeding with the usual classification.
We hypothesize translating or transliterating an under-resourced representation to a higher-resourced representation should help mitigate the problem of processing under-represented texts, thereby mitigating the long-tail distribution problem. Furthermore, LLMs are relatively better at simpler seq2seq tasks such as transliterating or translating.
We mix translation-based prompting with Non-Explanatory, Explanatory, and CoT-based prompting techniques.